OpenStudy (anonymous):

7x7+=49 and 2x8=16. .

OpenStudy (solomonzelman):

Okay, \(\LARGE\color{black}{ \frac{7}{8} - \frac{2}{7} }\) can you tell me the common denominator?

OpenStudy (anonymous):

First, you have to find the least common multiple of 8 and 7.

OpenStudy (anonymous):

@kalem03 What is the smallest number 8 and 7 go into?

OpenStudy (anonymous):

the least common multiple is 56?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

So what number do you multiply to 8 to get 56?

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{7}{8} - \frac{2}{7} }\) \(\LARGE\color{black}{ \frac{7\color{red}{\times 7}}{8\color{red}{\times 7}} - \frac{2\color{red}{\times 8}}{7\color{red}{\times 8}} }\)

OpenStudy (anonymous):

o.o

OpenStudy (solomonzelman):

notice, I am not changing any values. I am just multiplying top and bottom of each fraction. \(\LARGE\color{black}{ \frac{7\color{red}{\times 7}}{8\color{red}{\times 7}} - \frac{2\color{red}{\times 8}}{7\color{red}{\times 8}} }\) (can you simplify this expression, like to multiply all the red numbers?)

OpenStudy (anonymous):

Yeah, do what @SolomonZelman told you to do.

OpenStudy (anonymous):

Since 8 x 7 = 56, you have to multiply 7 to the numerator too, and since 7 x 8 = 56, you have to multiply 8 to 2 also.

OpenStudy (solomonzelman):

If you need any more help, reply....